A simple random sample of 5k race times for 32 competitive male runners aged 25-34 years old resulted in a mean 5k race time of 16.79 minutes.

The simple linear regression equation that fit the sample data was obtained. The regression equation was found to be ?' = 21.506 + 0.276? where ? represents the age of the runner in years and ? represents the 5k race time for a competitive male runner.

  1. What is the best predicted 5k race time for a competitive male runner 65 years of age?
Answer #1

Ans:

Given that

Regression equation:

y'=21.506+0.276 x

when x=65

Predicted race time,y'=21.506+0.276*65=39.45 minutes